OpenStudy (anonymous):

x-2y+3z=7 2x+y+z=4 -3X+2y-2z=-10

OpenStudy (anonymous):

are you like asking for an answer for each one of them?

OpenStudy (anonymous):

Write the initial augmented matrix for the system. Put the matrix into row echelon form. find the solution for x, y and z. x-2y+3z=7 2x+y+z=4 -3X+2y-2z=-10

OpenStudy (anonymous):

x-2y+3z=7 2x+y+z=4 -3X+2y-2z=-10